You want to buy a new surfboard 12 years from now. You have $600 in the bank today. You can earn 7% on your savings. If the surfboard will cost $5,000, how much will you have to save every year to have enough to buy it? Round your answer to 2 decimals, for example 100.12
Answer #1
Step 1 - First we have to find out the future value of $600 at the end of 12th year at 7% interest rate using future value of sum formula.
Step 2 - Then deduct this future value from $5000 ,the resulting amount will be the future value of yearly savings required.
Step 3 - We can find out this yearly savings using Future value of annuity formula.
Step 1
Future value of sum = P * (1+r)^n
P = amount in bank today = $600
r = interest rate per year = 7%
n = number of years = 12
Future value of sum = 600 * (1+0.07)^12 = $1,351.31
Step 2
Future value of yearly savings required = $5000 - $1351.31 = $3,648.69
Step 3
Future value of annuity = P * {[(1+r)^n -1]/r}
Future value of annuity = value calculated in step 2 = $3648.69
P = Yearly savings required to buy a new surfboard = ?
r = interest rate per year = 7%
n = number of years = 12
$3648.69 = P * {[(1+0.07)^12 -1]/0.07}
$3648.69 = P * 17.88845
P = 203.97
You have to save $203.97 every year to have enough money at the end 12th year to buy a surfboard.